OpenStudy (i_love_you):

Can anyone help me explain this please. A square field has an area of 479ft^2 . What is the approximate length of a side of the field? Give your answer to the nearest foot. I have to explain this in a informative way. Help please, thanks!

OpenStudy (boldjon):

A = 479 sq ft 479 = x^2 square root of 479 = square root of x^2 21.89... = x Rounded = 22 ft

OpenStudy (i_love_you):

Thank you :)

OpenStudy (boldjon):

np